OpenStudy (anonymous):

The diagonal of a a rectangle is 26m and its perimeter is 68m.what are its dimensions?

Parth (parthkohli):

The dimensions are given by length x width.

Parth (parthkohli):

Hint: Use some information to find the length and width.

OpenStudy (amistre64):

2a+2b=68 a^2 + b^2 = 26^2 might be a start

Parth (parthkohli):

\[ Perimeter = 2(l + w)\]First hint.

OpenStudy (anonymous):